The charges Q and q are placed at each of the diagonally opposite corners of a square . If the net electrical force on Q is 0 then find the ratio Q/q.

Grade:12

1 Answers

Arun
5 years ago

(KQ)/(2a2) = (√2)KQ/a2

|Q/q| = 2√2

So,

Q/q = -2√2

 

Hope it helps

 

Regards Arun

(AskIITians forum expert)